OpenStudy (hellomynameiskam):

Find the sum of 5/6 and 1/8. A. 6/11 B. 1 1/24 C. 23/24 D. 5/48

pooja195 (pooja195):

\[\huge~\rm~\frac{ 5 }{ 6 }+\frac{ 1 }{ 8 }=? \]

pooja195 (pooja195):

First let's start by finding an LCD Do you know what number we can use?

OpenStudy (hellomynameiskam):

Kinda

pooja195 (pooja195):

Tell me what you think :)

OpenStudy (hellomynameiskam):

8 ?

pooja195 (pooja195):

If we multiplied both of the denominators by 8 would they be the same?

pooja195 (pooja195):

They wouldn't... What number can we multiply 6 and 8 by to get the same denominator?

OpenStudy (hellomynameiskam):

would it be 6 then ?

pooja195 (pooja195):

6 x 6=36 8 x 6= 48 so no its not 6

pooja195 (pooja195):

Lets go down the list we know its not 1 lets try 2 8 x 2=? 6 x 2=?

OpenStudy (hellomynameiskam):

2 wouldnt work

pooja195 (pooja195):

Good And actually they dont need to be the same number we just need a number for both to make the numbers (denominators the same) 8 x 3=? 6 x 4=?

OpenStudy (hellomynameiskam):

it would be 24

pooja195 (pooja195):

Good :) \[\huge~\rm~\frac{ (4)5 }{ (4)6 }+\frac{ (3)1 }{ (3)8 }=?\]

pooja195 (pooja195):

Good :) \[\huge~\rm~\frac{ 20+3 }{ 24 }=?\] What would the answer be?

OpenStudy (hellomynameiskam):

23/24

pooja195 (pooja195):

Well done :-)